# Who Are We Missing?: A Principled Approach to Characterizing the Underrepresented Population

### v2

@@ 5 Learning $w$ via Functional Optimization
Consider any arbitrary dataset $\mathcal{A}_{n}=\{\mathbf{a}_{1},\dots,\mathbf{a}_{n}\}$, where $\mathbf{a}_{i}=(\mathbf{b}_{i},\mathbf{v}_{i})$ comprises a $p$-dimensional real-valued covariate vector $\mathbf{b}_{i}$ and a $k$-dimensional vector $\mathbf{v}_{i}$, and an objective function $\mathcal{L}$ that operates on a dataset $\mathcal{A}_{n}$. Additionally, consider a binary function $w:\mathbb{R}^{p}\to\{0,1\}$, such that $\forall(\mathcal{A}_{n},w)$, $\mathcal{L}(\mathcal{A}_{n},w)\geq 0$. Our goal is to determine an optimal function $w$ that minimizes $\mathcal{L}$:
$w^{*}\in\arg\min_{w}\mathcal{L}(\mathcal{A}_{n},w)$
where $w^{*}$ is the optimal but unknown function, and our additional objective is to interpretably characterize $w^{*}$.
@@ 5.1 Rashomon set of Optimal Trees (ROOT)
Given the stochastic nature of our tree construction method, there is no assurance that a single tree is optimally minimizing the objective function. To address this challenge, we generate a set of $M$ trees, with a higher $M$ more desirable. For practical considerations, we construct approximately $M=5000$ trees in each of our empirical studies, both with synthetic and real data. We then rank these trees based on their respective objective function values and select the top $m$ trees with the smallest objective values. We shape our final $w$ by aggregating predictions from each tree using a maximal voting approach. To ensure interpretability, we keep the value of $m$ relatively small, typically ranging between 5 and 15. We call this ensemble of selected trees the Rashomon set Of Optimal Trees (ROOT) [^1] Rashomon is a classic Japanese movie in which an event is given non-degenerate interpretations by the individuals involved. It provides different and sometimes contradictory points of view on the same event. Here, each of the trees in the set provides almost equally good but the different characterization of underrepresented subpopulations – thus we refer to this set as a Rashomon Set. Analogously, there is also literature in Interpretable Machine Learning that focuses on understanding the properties and behavior of almost equally good but different models (Breiman,, 2001; Semenova et al.,, 2022).. Although each tree in ROOT is nearly equally effective, they provide slightly different interpretations. Leveraging the sparsity of these trees and the manageable size of ROOT, we further create a concise (albeit non-optimal) single-tree explanation for the $w$ characterized by ROOT. This single-tree explainer learns a sparse representation that mimics the prediction function learned by ROOT – we also refer to this as the characteristic tree. We delineate the algorithm in Appendix B.
@@
and $n_{1}=\sum_{i=1}^{n}S_{i}.$
6 Synthetic Data Experiments
Now, we study the performance of our proposed approach via synthetic data experiments. In particular, we are interested in understanding the extent to which our approach (i) produces interpretable characterizations of underrepresented populations, (ii) reduces the variance of the WTATE relative to the variance of the TATE, and (iii) handles high-dimensional and non-linear data in three data-generative processes (DGPs) with different features.
6.1 Data Generative Process
These three synthetic DGPs are designed to be akin to different real-world scenarios and illustrate important features of our approach. We call them (i) Community, (ii) Box, and (iii) High-dimensional DGPs due to their respective properties.
Community DGP.
$P(C_{i}=A)=\frac{3}{4}\textrm{ and }P(C_{i}=B)=\frac{1}{4}$
Box DGP.
$X_{i,0},X_{i,1},\ldots,X_{i,100}\sim\mathcal{U}(0,1).$
$Y_{i}(1)=Y_{i}(0)+\log(Y_{i}(0)+1).$
In summary, the Box DGP model involves high-dimensional covariates, non-linear response functions, and varying trial participation probabilities based on covariate characteristics.
High-dimensional DGP.
Similar to the Box DGP, here too we consider a high-dimensional setup with 100 iid covariates per individual: $\{X_{i,0}\dots X_{i,100}\}\overset{iid}{\sim}\mathcal{N}(0,1)$. Unlike the Box DGP, all 100 covariates contribute towards an individual’s likelihood of participating in the trial: $P(S_{i}=1\mid\mathbf{X}_{i})=\mathrm{expit}(\alpha_{0}^{T}\mathbf{X}_{i})$. The control potential outcome and the treatment effect are linear functions of covariates: $Y_{i}(0)=\beta_{0}^{T}\mathbf{X}_{i}+\epsilon_{i}$, and $Y_{i}(1)=Y_{i}(0)+\beta_{1}^{T}\mathbf{X}_{i}+\epsilon_{i}$ where $\epsilon_{i}\sim\mathcal{N}(0,1)$. Here, the parameters $\alpha_{0}$, $\beta_{0}$, and $\beta_{1}$ are also drawn from a standard normal distribution.
6.2 Analysis
We first examine the extent to which our approach reduces the variance of the WTATE estimates compared to the TATE estimates. Table 1 shows the empirical bias as well as standard errors for each of the four approaches as well as the estimates on the unpruned (or unrefined) target. The precision measured as the standard error is a function of study data as well as the estimator. We use the same IPW WTATE estimator across all 4 approaches, however, empirical bias might vary given the underlying data sample. Focusing on the standard errors (precision), we find that for both Community and Box DGPs, the ROOT approach produces the most precise estimates. For High-dimensional DGP, however, the Linear-w approach tends to outperform other approaches. This is primarily because all the relationships in this DGP are linear and thus $w(\mathbf{X})=\gamma^{T}\mathbf{X}$ is the “correct” specification. It is worth noting that ROOT is the second-best approach in this scenario. Thus, our experimental results indicate that our approach improves the precision of the TATE estimate without compromising the bias. Further, given the nonlinearity of the Box DGP, and the high-dimensional nature of both the Box and High-dimensional DGPs, our results also support the ability of our approach to handle a complex and high-dimensional dataset.
Method
Original
Target
Estimate
ROOT 1-Tree Linear Indicator
Community DGP
Bias -5.130 1.500 -5.581 -3.410 -6.802
Std. Error 3.073 0.449 0.985 2.995 0.770
Box DGP
Bias 0.220 -0.002 -1.088 0.220 0.091
Std. Error 0.910 0.324 0.409 0.910 0.763
High-dim. DGP
Bias -0.310 -0.126 -0.190 -0.342 -0.392
Std. Error 0.191 0.162 0.169 0.127 0.170
Table 1: Table of Precision (measured as standard error) and Accuracy (measured as empirical bias) for the three DGPs described in Section 6.1, and four approaches described in Section 5. ROOT is our proposed approach while the other three are baseline comparison methods.
Next, we discuss how one can use the ROOT approach to produce interpretable characteristics of the underrepresented groups. In Figure 1, we show three trees characterizing the refined study population for the three DGPs. Further, Figure 2 shows the data distribution color-coded based on the predicted values of $w(\mathbf{X})$ using ROOT for each unit in the dataset across the two most important features for each of the DGPs that are both outcome modifiers as well as the sample selectors. Our results shown in these figures indicate that ROOT is not only correctly able to identify the underrepresented groups or communities but also produces a tree that can be used to communicate their characteristics easily.
Unlike the results for the Community and the Box DGPs, in the case of the High-dimensional DGP, the decision boundary in the two-dimensional projection is not sharp. This is because in this scenario all of the 100 covariates are moderators and relate to trial participation and no two-dimensional slicing of the covariate space can perfectly show the decision boundary. Further, the characteristic tree is also significantly deeper in the High-dimensional DGP. This is because of the large number of moderators being almost equally important, there is no shallow tree that can accurately characterize the groups.
(a) Community DGP
(b) Box DGP
(c) High-dim. DGP
Figure 1: Tree characterizing indicating the study population in Blue and the underrepresented subgroups in Orange for all three DGPs
Refer to caption Refer to caption Refer to caption
(a) Community DGP (b) Box DGP (c) High-dim. DGP
7 Conclusion
@@
Our paper addresses the common issue of underrepresented populations in clinical trials when generalizing treatment effects to a target population. The presence of underrepresented subgroups can significantly undermine the accuracy and precision of estimated target treatment effects, thereby impacting subsequent decision-making processes. In response to this challenge, we propose a principled approach to characterize a refined target population, optimizing for precision systematically. This facilitates a clear understanding of subpopulations for which the trial inferences cannot be extended with high precision. As a potential future direction, it can empower the design of trials with a specific focus on these underrepresented subgroups.
@@
Our experiments with synthetic data demonstrate that our approach enhances the precision of estimates compared to alternative methods while maintaining interpretability. Furthermore, we apply our method to extend inferences from the START trial, generalizing the comparative treatment effect of methadone vs. buprenorphine on OUD relapse within 24 weeks after randomized assignment. Our analysis identifies underrepresented subgroups in the trial using a shallow, easily interpretable tree. Post-pruning, our refined analysis significantly improves the precision of the generalized effect estimate.
@@
While our approach significantly enhances the quality of analyses and subsequent decision-making, it also highlights open questions. For example, similar to Crump et al., (2009); Li et al., (2019), we use the data to define the estimand, which has implications for inference in this context. We conceptualize our approach as a two-step procedure, encompassing design and analysis. In the design step, we leverage ROOT alongside trial data to refine the target population. In the subsequent analysis step, we consider this redefined estimand fixed and proceed to estimate the target average treatment effect. An open question pertains to whether accounting for uncertainty in estimating the weights $w$ is necessary and how it impacts TATE estimation and inference.
Additionally, our method assumes positivity (A.1.) and does not focus on identifying completely unrepresented populations but rather on identifying populations that are somewhat but not sufficiently represented.
@@
Notably, while ROOT is employed for estimating $w$, it is an algorithm that uses randomization, and its general optimality is not guaranteed. ROOT’s proposed solution achieves optimality as the number of grown trees approaches infinity, representing an exhaustive exploration of the entire tree space. However, for a finite set of trees, the optimality gap remains unknown. Future work will concentrate on exploring the computational and algorithmic properties of ROOT, aiming to quantify this optimality gap.
While ROOT is tailored for characterizing target populations and minimizing variance, its versatility extends beyond the presented application. Our ongoing work will focus on extending ROOT to a broader setting and adapting it for non-binary $w$ functions.

### v4

@@ 5 Learning $w$ via Functional Optimization
Consider any arbitrary dataset $\mathcal{A}_{n}=\{\mathbf{a}_{1},\dots,\mathbf{a}_{n}\}$, where $\mathbf{a}_{i}=(\mathbf{b}_{i},\mathbf{v}_{i})$ comprises a $p$-dimensional real-valued covariate vector $\mathbf{b}_{i}$ and a $k$-dimensional vector $\mathbf{v}_{i}$, and an objective function $\mathcal{L}$ that operates on a dataset $\mathcal{A}_{n}$. Additionally, consider a binary function $w:\mathbb{R}^{p}\to\{0,1\}$, such that $\forall(\mathcal{A}_{n},w)$, $\mathcal{L}(\mathcal{A}_{n},w)\geq 0$. Our goal is to determine an optimal function $w$ that minimizes $\mathcal{L}$, $w^{*}\in\arg\min_{w}\mathcal{L}(\mathcal{A}_{n},w)$ where $w^{*}$ is the optimal but unknown function, and our additional objective is to interpretably characterize $w^{*}$.
@@ 5.1 Rashomon set of Optimal Trees (ROOT)
Given the stochastic nature of our tree construction method, there is no assurance that a single tree is optimally minimizing the objective function. To address this challenge, we generate a set of $M$ trees, with a higher $M$ more desirable. For practical considerations, we construct approximately $M=5000$ trees in each of our empirical studies, both with synthetic and real data. We then rank these trees based on their respective objective function values and select the top $m$ trees with the smallest objective values. We shape our final $w$ by aggregating predictions from each tree using a maximal voting approach. To ensure interpretability, we keep the value of $m$ relatively small, typically ranging between 5 and 15. We call this ensemble of selected trees the Rashomon set Of Optimal Trees (ROOT). Although each tree in ROOT is nearly equally effective, they provide slightly different interpretations. Leveraging the sparsity of these trees and the manageable size of ROOT, we further create a concise (albeit non-optimal) single-tree explanation for the $w$ characterized by ROOT. This single-tree explainer learns a sparse representation that mimics the prediction function learned by ROOT – we also refer to this as the characteristic tree. We delineate the algorithm in Appendix B.
@@
and $n_{1}=\sum_{i=1}^{n}S_{i}.$ o
6 Empirical Results
6.1 Summary of Synthetic Data Analyses
We conducted synthetic data experiments to evaluate the effectiveness of our proposed approach, ROOT, in generating interpretable representations while accurately characterizing underrepresented groups. These experiments were designed to assess ROOT’s performance in high-dimensional and non-linear settings, comparing it to alternative methods. We employed four distinct data generative procedures, including one that closely mimics real-world data on medications for opioid use disorder. Detailed descriptions of the data generative procedures and the subsequent analyses are provided in Appendix D. As shown in Table 2, refining the study population using ROOT led to significant improvements in precision. Additionally, Figure 5 illustrates that ROOT produces interpretable results in the form of sparse or shallow decision trees. However, while ROOT performs optimally across most data generating processes (DGPs), it struggles to maintain sparsity when the effect modifiers themselves are high-dimensional. This limitation highlights a potential area for future refinement.
6.2 Insufficiently Represented Groups in MOUD Trials
Now, we apply our methodology to the medication for opioid use (MOUD) data where we are interested in transporting the treatment effect from the Starting Treatment With Agonist Replacement Therapies (START) trial to the population of individuals in the US seeking treatment for opioid use disorder, using the Treatment Episode Dataset-Admissions 2015-2017. We are interested in (i) the characteristics of the subpopulation for which we can precisely estimate the TATE using the trial evidence, (ii) the TATE estimate for this subpopulation, and (iii) the characteristics identifying those who are treated in TEDS-A but underrepresented in the trial cohort.
6.2.1 Data Description
We first begin the discussion by describing the two datasets used: the trial and population data.
Trial Data.
The Starting Treatment With Agonist Replacement Therapies (START) trial, initiated in April 2006, was a multi-center study designed to compare the effects of buprenorphine and methadone (Saxon et al.,, 2013; Hser et al.,, 2014). It involved 1271 participants who were randomly assigned to either the buprenorphine or methadone arms in a 2:1 ratio. We chose this trial as it is one of the largest MOUD trials conducted as part of the NIDA Clinical Trials Network, which are trials that are conducted in real-world treatment settings, designed to improve generalizability (Tai et al.,, 2011). The trial’s findings indicated that methadone was associated with better retention in treating opioid use disorder compared to buprenorphine (Hser et al.,, 2014). However, it must be noted that the trial compares methadone to an atypical buprenorphine administration, deviating from real-world practices. Individuals had to return to the study site daily for distribution of buprenorphine. In usual care settings, buprenorphine is distributed on a 2-6 week prescription interval .
In our analysis, the outcome is relapse by 24 weeks following the treatment assignment. Relapse is delineated as the consistent use of non-study opioids at least once per week for four consecutive weeks or daily usage of non-study opioids for seven continuous days. Data on opioid use is collected through urine drug screens and self-reported during follow-up visits. The assessment of relapse initiates 20 days post-randomization to account for positive urine screens that may be attributed to opioid use during medically managed withdrawal or residual drug presence during stabilized treatment.
Target Sample.
We employed the Treatment Episode Dataset - Admission (TEDS-A) dataset from 2015-2017 as a representative cross-section of the target demographic. This dataset includes individuals who sought assistance from publicly funded substance use treatment programs, spanning inpatient, outpatient, and residential facilities under the management of the Substance Abuse and Mental Health Services Administration. The TEDS-A dataset utilized in this study comprised information from 48 states (excluding Oregon and Georgia) and the District of Columbia. Notably, data for South Carolina was unavailable for the year 2015. Our analysis focuses on a cohort of 740,015 individuals who fulfilled two specific criteria: (i) the inclusion of Medication-Assisted Treatment (MOUD) in their treatment plan and (ii) a DSM-IV-TR diagnosis of opioid dependence or opioid abuse, or individuals entering treatment with documented heroin use, non-prescription methadone use, or other/synthetic opioid use.
Data Summary.
For our analysis, we consider methadone as $T=1$ and buprenorphine as $T=0$. For the outcome, $Y=1$ indicates relapse by 24 weeks after treatment assignment. We only consider the pre-treatment covariates observed in both the trial and target data. Our set of pre-treatment covariates includes the participants’ age, race and biological sex along with substance use history (amphetamine, benzodiazepines, cannabis and IV drug use) assessed at the initiation of MOUD treatment. See Table 3 for a listing of all of the covariates included.
6.2.2 Analysis
Trial ATE.
We first estimate the average treatment effect in the trial sample. Individuals assigned to methadone exhibit a $10.04$ percentage point reduction in the likelihood of relapse by 24 weeks compared to those on buprenorphine i.e. the experimental ATE is 10.0410.04-10.04- 10.04222This particular comparative treatment effect compares methadone to an atypical administration of buprenorphine in which buprenorphine is administered during daily or near-daily clinic visits to match the administration of methadone, but which is markedly different from administration in the real world in which buprenorphine is prescribed every 2-6 weeks. Therefore, this particular comparison involves a much more burdensome administration of buprenorphine., with a standard error of $3.66$.
Estimating TATE.
Next, we employ the IPW estimator to extend the inferences from the START trial to the population of opioid users represented by the TEDS-A dataset. This relies on assumptions A.1.-A.4. We discuss the validity of these assumptions in Appendix F. The TATE estimate using IPW indicates a modest advantage for methadone, with a $-9.72$ percentage points reduction in relapse likelihood, albeit with a wide standard error of $4.680$. Note that the estimated TATE has a relatively larger standard error than the trial sample average treatment effect.
Shifting the Goalpost.
Now, we focus on refining the analysis by employing a two-stage approach: (i) the design stage identifies the underrepresented subpopulations and subsequently excludes them from the study population, and (ii) the analysis stage estimates the TATE for this refined study population. For the design stage: we use four different approaches including our proposed one using ROOT, two traditional approaches using selection scores ($\ell(\mathbf{x})$) and using “indicator” $w(\mathbf{x})$, and compare the results using each of them. For the analysis stage: we use the IPW estimator to estimate TATE using the refined target population. We compare the results via ROOT with the results using the selection score.
Figure 1: Swarm-plot and violin-plot showing the distribution of the normalized selection odds $\ell(\mathbf{x})/\ell$ for the units in the trial sample ($S=1$) characterized as $w(\mathbf{x})=0$ (i.e. “underrepresented”) and $w(\mathbf{x})=1$ using (a) selection odds with predefined threshold of $0.87$ on $\ell(\mathbf{x})/\ell$, (b) selection odds with the optimal threshold for $\ell(\mathbf{x})/\ell$ and (c) using ROOT.
The traditional selection score-based approaches are akin to an approach proposed in Crump et al., (2009) and Li et al., (2019) to identify underrepresented populations using $\ell(\mathbf{x})$, and in particular the normalized selection odds, i.e., $\ell(\mathbf{x})/\ell$. In this approach, the problem reduces to finding a threshold $\ell_{\spadesuit}$ such that units with $\ell(\mathbf{x})/\ell$ less than $\ell_{\spadesuit}$ are underrepresented. We consider two versions of such a selection score-based “design-stage”: the first version uses a pre-defined $\ell_{\spadesuit}=0.87$ (we select this using the natural breakpoint in the distribution of normalized selection odds), and the second version learns the optimal $\ell_{\spadesuit}$ through an “indicator” $w(\mathbf{x})$ approach, which chooses $w(\mathbf{x})=\sum_{i}w_{i}\mathbf{1}[\mathbf{x}_{i}=\mathbf{x}]$ and optimizes for $\{w_{i}\}$. Lastly, we use our proposed approach, ROOT, to construct $w(\mathbf{x})$ using trees.
Table 1 shows the WTATE point estimate and standard errors for each of the four design stage strategies along with the original TATE estimates and sample ATE. Figure 6 shows the distribution, along $\ell(\mathbf{x})/\ell$, of the refined target population and the subpopulations characterized as underrepresented for each of the four approaches. The “indicator” $w(\mathbf{x})$ yields the WTATE with the smallest standard error followed by the one using ROOT. However, “indicator” $w(\mathbf{x})$ results in pruning many more units from the target population compared to ROOT, thus limiting the population for inference. Further, “indicator” $w(\mathbf{x})$ also characterizes a significant subpopulation with $\ell(\mathbf{x})/\ell$ greater than 1 as underrepresented – this is potentially a sign of overfitting. Overall, ROOT produces smaller standard errors compared to both the selection score-based approaches (third column of Table 1) while also ensuring that the refined target population is as large as possible (last column of Table 1), avoiding overfitting (last column of Table 1) and ensuring interpretability (Figure 2).
Table 1: Table of estimands corresponding point estimate, standard error, and relative size of the refined target population measured via $\mathbb{E}[w(\mathbf{X})]$ for the four choices of design stage strategies.
Estimand Point Est Std Err Remarks $\mathbb{E}[w(\mathbf{X})]$
Trial ATE -10.040 3.666 on START trial sample -
TATE -9.724 4.680 using full TEDS-A data 1.000
WTATE (selection score) -8.641 4.750 predefined threshold 0.884
WTATE (selection score) -10.171 4.676 optimized threshold 0.938
WTATE (Indicator) -27.744 4.612 - 0.888
WTATE (ROOT) -12.182 4.666 - 0.993
Characterizing Underrepresented Populations.
One key aspect of ROOT’s design stage is its interpretability along with flexibility, which allows for communication of the characteristics of the underrepresented population that is excluded from the refined study population. Figure 2 presents a decision tree that characterizes these subgroups. We identify that the following two subgroups are underrepresented in the START trial: (i) Black people with a history of amphetamine use but no history of cannabis use, and (ii) Hispanic women with a history of amphetamine and benzodiazepine use.
Figure 2: Decision Tree characterizing the underrepresented population. The orange nodes and leaves indicate underrepresented subgroups that can be excluded from the TATE analysis for better precision.
One key difference between these insufficiently represented groups and the remaining groups is the distribution of conditional average treatment effects (as shown in Figure 3(a). Figure 3(b) performs a deeper dive into understanding the distribution of conditional average treatment effects as a function of covariates. Our results indicate that while for the majority of the population methadone (T=1) appears to be the superior treatment, for the following two subgroups buprenorphine (T=0) is the superior treatment: individuals who do not have a history of cannabis use and are either (i) racially Black or (ii) biological males. Of these two groups, the average normalized selection odds, $\ell(\mathbf{x})/\ell$, and corresponding CATE estimates for racially Black individuals is $0.909$ and $+26.770$, and for the biological males is $1.001$ and $+19.978$, respectively. The heterogeneity of CATE along with the selection odds less than one implies a potential lack representation on effect modifiers. These analysis results are in agreement with the ROOT-based characteristic tree which identifies Black individuals without cannabis use history as insufficiently represented in the START trial.
Refer to caption Refer to caption
(a) (b)
Figure 3: (a) Distribution of conditional average treatment effects for units with $w(\mathbf{X}_{i})=1$ (well represented) and $w(\mathbf{X}_{i})=0$ (characterized underrepresented) via ROOT. (b) Tree explaining the distribution of conditional average treatment effects across various subgroups. Here, the yellow leaves indicate methadone is more effective compared to buprenorphine while the green leaves indicate buprenorphine is more effective than methadone in that subpopulation.
7 Discussion & Conclusion
@@
Our paper addresses the common issue of insufficiently represented (also referred here as underrepresented) populations in clinical trials when generalizing treatment effects to a target population. The presence of insufficiently represented subgroups can significantly undermine the accuracy and precision of estimated target treatment effects, thereby impacting subsequent decision-making processes. In response to this challenge, we propose a principled approach to characterize the underrepresented groups by optimizing for precision systematically. This facilitates a clear understanding of subpopulations for which precise estimation of treatment effect is difficult. Such an exercise of interpretable characterization is of value for but not limited to (i) understanding contributions to research-to-practice gaps, such as in medication for opioid use disorder studies, and (ii) designing future studies to understand effects for the underrepresented groups.
@@
On Refined Estimand.
Our procedure redefines the estimand to focus on a causal quantity within a well-defined subgroup that has sufficient support in the data. The primary goal is to identify and characterize underrepresented regions of the covariate space in the trial cohort. Once these regions are identified, they are excluded from our analysis, which means dropping units in these regions from both the experimental and target samples. Consequently, the refined population is likely not identical to the original trial cohort. This approach involves a fundamental tradeoff: while our generalized estimates for the refined estimand, WTATE, have greater credibility, they are less general than the original estimand, TATE. Note that while it is crucial to ensure balance on effect moderators when generalizing or transporting inferences, extrapolating in the direction of features that are not effect modifiers is feasible without any loss of precision.
On Interpretability.
Sparse decision trees are learned across the literature for their ability to produce interpretable results. However, there are possibly other decision trees that are nearly as optimal as the learned tree and have an alternate interpretation (Xin et al.,, 2022; Semenova et al.,, 2022). Thus, analogous to the traditional interpretable ML literature (Xin et al.,, 2022; Ruggieri,, 2019), we use the Rashomon set of the optimal trees as the interpretations are more comprehensive and stable. Here, the Rashomon set is constrained to contain the $k$ best trees with the smallest objective value out of $M$ generated trees (e.g., in our case study we choose $k=5$). The Rashomon set itself is considered “interpretable” because it is small and each element is individually interpretable. However, some researchers may desire a simpler characterization. Thus, we also summarize (aggregate) the representations produced by the trees in the Rashomon set via a characteristic tree, providing a result that is both interpretable and simple, though containing less expressive power than the full Rashomon set of optimal trees. Our software package outputs both this characteristic tree and the full Rashomon set with the respective objective values for each tree.
Takeaways from Empirical Study.
@@
Our synthetic data experiments demonstrate that in scenarios where the number of moderators is relatively small, our approach enhances the precision of estimates compared to alternative methods while maintaining interpretability. We also explore the limits of our proposed approach via a high-dimensional scenario with a large number of moderators where it is challenging to construct a sparse representation of the underrepresented population. Lastly, we also compare the performance of ROOT and other baseline approaches on a synthetic DGP that mimics the distribution of our real-world MOUD case study data. Here, ROOT outperforms other methods as well as yields an interpretable representation of the underrepresented population that is in congruence with the DGP. Furthermore, we apply our method to extend inferences from the START trial, generalizing the comparative treatment effect of methadone vs. buprenorphine on OUD relapse within 24 weeks after randomized assignment. Our analysis identifies underrepresented subgroups in the trial using a shallow, easily interpretable tree. Post-pruning, our refined analysis significantly improves the precision of the generalized effect estimate.
On Data-Adaptive Estimand.
@@
While our approach significantly enhances the quality of analyses and subsequent decision-making, it also highlights open questions. For example, similar to Crump et al., (2009); Li et al., (2019), we use the data to define the estimand, which has implications for inference. We conceptualize our approach as a two-step procedure, encompassing design and analysis. In the design step, we leverage ROOT alongside trial data to refine the target population. In the subsequent analysis step, we consider this redefined estimand fixed and proceed to estimate the target average treatment effect. An open question pertains to whether accounting for uncertainty in estimating the weights $w$ is necessary and how it impacts TATE estimation and inference.
On Positivity Violation
For consistent estimation of the target average treatment effect, representativeness is required only on effect moderators that differ in distribution between the trial and target, rather than on every feature. However, in many real-world scenarios, we do not know which features act as effect moderators. Our procedure leverages observed data to implicitly identify these effect moderators. In this context, our method assumes positivity and focuses on identifying “underrepresented” populations – those that are present but not sufficiently so – rather than completely “unrepresented” populations. To illustrate this with an example, consider a binary feature encoding “pregnancy”: if the trial includes only non-pregnant individuals and excludes pregnant individuals, we cannot ascertain if pregnancy status is an effect modifier using the observed data alone. Consequently, we cannot determine the necessity of balancing pregnancy status or if pregnant individuals are underrepresented. Thus, we recommend preprocessing the data to a priori prune regions with positivity violations (i.e., $P(S=1\mid\mathbf{X})=0$) based on domain knowledge or trial admission/exclusion criteria.
On Existence of Unobserved Confounder.
In addition to positivity, generalizability or transportability also depends on the external validity of RCT (A3) — this assumption is also referred to as S-admissibility or sample exchangeability. There is an extensive literature that focuses on the partial identification of transported/generalized treatment effects and sensitivity analysis of the estimation step to violations of the external validity assumption (Dahabreh et al.,, 2022; Colnet et al.,, 2021; Nguyen et al.,, 2017; Huang,, 2024; Zeng et al.,, 2023; Nie et al.,, 2021). However, to the best of our knowledge, no prior work has discussed the sensitivity of a data-adaptive estimand, like the one we propose here, to a violation of the external validity assumption. This is a key gap, as mismeasurement or the absence of an important effect moderators can change the eventual estimand of interest. Future work could focus on developing sensitivity analysis tools for data-adaptive estimands such as that proposed here.
On Optimality Guarantee of ROOT.
@@
Furthermore, while ROOT is employed for estimating $w$, it is an algorithm that uses randomization, and its general optimality is not guaranteed. In Appendix C, we show that ROOT’s proposed solution achieves optimality, given a fixed dataset, as the number of grown trees approaches infinity, representing an exhaustive exploration of the entire tree space. However, for a finite set of trees, the optimality gap remains unknown. Future work will concentrate on exploring the computational and algorithmic properties of ROOT, aiming to quantify this optimality gap and quantify sampling uncertainty. While ROOT is tailored for characterizing target populations and minimizing variance, its versatility extends beyond the presented application. Our ongoing work will focus on extending ROOT to a broader setting and adapting it for non-binary $w$ functions. Further, we will explore the convergence, stabilization and finite-sample optimality gap of ROOT-based optimization.